OpenStudy (roadjester):

\(\Psi = \sqrt{\dfrac 2 L} sin{(\dfrac {2\pi x}{L})}\) for 0\(\le\)x\(\le\) L. Determine the expectation value. The expectation value is calculated as \[ = \int_{-\infty}^{+\infty} \Psi^* x \Psi dx\] where \(\Psi^*\) is the complex conjugate of \(\Psi\). Anyone know how to do the calculation? (a) Determine the expectation value of x. (b) Determine the probability of finding the particle near \(\dfrac 1 2\)L by calculating the probability that the particle lies in the range 0.490L\(\le\) x\(\le\) 0.510L. (c) What If? Determine the probability of finding the particle near \(\dfrac 1 4\)L by calculating the probability that the particle lies in the range 0.240L \(\le\) x \(\le\) 0.260L. (d) Argue that the result of part (a) does not contradict the results of parts (b) and (c).

OpenStudy (roadjester):

@ybarrap @agent0smith

OpenStudy (roadjester):

@niksva

OpenStudy (anonymous):

complex conjugate of wave function would be same so we will get \[\frac{ 2 }{ L } \int\limits_{0}^{L} x*\sin^2(\frac{ 2 pix }{ L })\] is that clear?

OpenStudy (roadjester):

But why is it the same? Aren't I working with complex numbers?

OpenStudy (anonymous):

yeah we are working with the complex number in this case we have the real wave function not an imaginary wave function

OpenStudy (roadjester):

How do you determine that? (Sorry, new to Quantum Physics)

OpenStudy (anonymous):

u must have done the derivation of wavefunction for 1D potential well??????

OpenStudy (roadjester):

sorry, computer crashed. Okay, because it's oscillating, it has to be a Sin/Cos wave. Due to boundary conditions, It has to be the sin wave.

OpenStudy (roadjester):

Is that what you're referring to?

OpenStudy (anonymous):

sorry owl was not allowing me to come here :-P

OpenStudy (roadjester):

Me too

OpenStudy (roadjester):

Anyways ^^^^^^^ was that what you were talking about?

OpenStudy (anonymous):

yeah because of the boundary condition we have sine wave

OpenStudy (anonymous):

now can u determine expectation value?

OpenStudy (roadjester):

Isn't it just integration?

OpenStudy (roadjester):

Is there more to it? (I just haven't done calculus in a while...=( )

OpenStudy (anonymous):

right now its just integration use double cosine angle formulae u will surely get the expectation value of x

OpenStudy (roadjester):

I'm going to take a little while.

OpenStudy (anonymous):

cool take your time

OpenStudy (roadjester):

\[sin^2(\theta /2)={\dfrac 1 2} (1-cos\theta)\] Is this the one you're talking about?

OpenStudy (anonymous):

yeah

OpenStudy (roadjester):

I'm not sure how to manipulate that since it's \(\theta /2\) on the inside.

OpenStudy (anonymous):

can we write sin^2(theta) = 1/2 (1-cos(theta) ) ???????

OpenStudy (roadjester):

I think I recognize that one.

OpenStudy (anonymous):

i m now writing all the steps in one go

OpenStudy (roadjester):

\(\huge \dfrac {1}{L} \int_{-\infty}^{+\infty} x[ 1-cos(2{\dfrac{2\pi x}{L}}) ] dx\)

OpenStudy (roadjester):

That's what I get so far

OpenStudy (anonymous):

\[\frac{ 2 }{ L }\int\limits_{0}^{L}x*\sin^2(n pix/L)\] \[\frac{ 2 }{ L }\int\limits_{0}^{L} x*\frac{ 1 }{ 2 }(1-\cos(\frac{ 4 pix }{ L })) dx\]

OpenStudy (roadjester):

Why 0 to L?

OpenStudy (anonymous):

Limits should be in between 0 to L as it is a potential box of dimension L

OpenStudy (roadjester):

Is this integration by parts? due to the x on the outside?

OpenStudy (anonymous):

\[\frac{ 1 }{ L }\int\limits_{0}^{L} \ (x- xcos(\frac{ 4 pix }{ L })) dx\] for xcos(4pix/L) go for integration by part while other one is simpler

OpenStudy (roadjester):

ok so I get \[\dfrac L 2-\dfrac 1 L\int_0^L {x cos({\dfrac {4\pi x}{L}})}dx\]

OpenStudy (anonymous):

yeah now solve the integral i feel the integral will come out to be zero as this is an odd function

OpenStudy (roadjester):

How do you figure it's odd?? The fact that it's a sine function?

OpenStudy (anonymous):

we have a function f(x) = xcos(4pix/L) function is odd when f(-x) = - f(x) in the above case condition f(-x) = -f(x) satisfies hence it is an odd function

OpenStudy (roadjester):

But its limits are from 0 to L, not -L to L

OpenStudy (anonymous):

yeah it is it was just my feeling :-) we will solve this out dont worry

OpenStudy (roadjester):

one sec; almost done with the integration by parts.

OpenStudy (roadjester):

Ok yeah, I think you were right

OpenStudy (roadjester):

This is my final answer

OpenStudy (roadjester):

\(\Large \dfrac L 2-\dfrac {x}{4\pi}sin({\dfrac {4\pi x}{L}})\)

OpenStudy (anonymous):

\[x*\frac{ \sin(\frac{ 4 \pi x }{ L } }{ \frac{ 4 \pi }{ L } }-\int\limits_{}^{}\frac{ d }{ dx }(x) \frac{ \sin \frac{ 4 \pi x }{ L } }{ \frac{ 4\pi }{ L } }\] \[x*\frac{ \sin \frac{ 4 \pi x }{ L } }{ \frac{ 4 \pi }{ L } } - \int\limits_{}^{}\frac{ L }{ 4 \pi }* \sin (\frac{ 4 \pi x }{ L})dx\] can u do it now?

OpenStudy (anonymous):

\[\frac{ Lx }{ 4\pi } \sin \frac{ 4 \pi x }{ L } +\frac{ L^2 }{ 16 \pi^2 } \cos(\frac{ 4 \pi x }{ L })\] now apply limit

OpenStudy (roadjester):

I already told you my final answer, but it had x in it...

OpenStudy (anonymous):

u did not apply limit right u have to put upper limit L and lower limit 0 in place of x

OpenStudy (anonymous):

my final answer is \[\frac{ L }{ 2 } +\frac{ L }{ 16 \pi^2 }\]

OpenStudy (roadjester):

\[\Large \dfrac L 2-\dfrac {x}{4\pi}sin({\dfrac {4\pi x}{L}})\] so if I apply limits 0 and L to this?

OpenStudy (roadjester):

I get L/2.... I'm so confused

OpenStudy (anonymous):

u have done integration by parts wrong see the steps done by me first

OpenStudy (roadjester):

\[\frac{ Lx }{ 4\pi } \sin \frac{ 4 \pi x }{ L } +\frac{ L^2 }{ 16 \pi^2 } \cos(\frac{ 4 \pi x }{ L })\] In the first term, why is there an L in the numerator? didn't the 1/L at the beginning cancel it out?

OpenStudy (roadjester):

\[\frac{ \color {red }{\Large L} x }{ 4\pi } \sin \frac{ 4 \pi x }{ L } +\frac{ L^2 }{ 16 \pi^2 } \cos(\frac{ 4 \pi x }{ L })\]

OpenStudy (roadjester):

@niksva I put the integral into wolfram alpha, it's L/2.

OpenStudy (roadjester):

I need help with parts b-d.

OpenStudy (roadjester):

@theEric

OpenStudy (anonymous):

@roadjester in part b and c integral will will be same only the limit has been changed in the first part we are expecting the particle to be find in one dimensional well of length L now in other parts , range of well has been reduced

OpenStudy (anonymous):

this is your solution of first part Last day I was just doing here on this forum today i took the paper and pen

OpenStudy (roadjester):

Thanks @niksva I managed to solve the rest of the problem but thanks for the help.